MARC (marc_d):

Two dice are rolled.What is the probability of getting the numbers whose sum is 5 or their difference is 1?

MARC (marc_d):

@ganeshie8

MARC (marc_d):

A=getting whose sum is 5 ={1,2,3,4} B=getting their difference is 1 ={1,2,3,4,5,6}

MARC (marc_d):

is correct?

ganeshie8 (ganeshie8):

A=getting whose sum is 5 ``` 1+4 2+3 3+2 4+1 ```

MARC (marc_d):

ok

ganeshie8 (ganeshie8):

Let's find the number of ways for B=getting their difference is 1

MARC (marc_d):

aite it would be.. 2-1 3-2 4-3 5-4 6-5

ganeshie8 (ganeshie8):

I think so

MARC (marc_d):

ok

ganeshie8 (ganeshie8):

What must we do next ?

OpenStudy (kropot72):

Pairs, with the difference being 1: 5,6 4,5 5,4 4,3 3,4 3,2 2,3 2,1 1,2

OpenStudy (kropot72):

and 6,5

ganeshie8 (ganeshie8):

Ahh nice, I think it makes more sense to include above

MARC (marc_d):

Yes :)

ganeshie8 (ganeshie8):

aite it would be.. 2-1 3-2 4-3 5-4 6-5

ganeshie8 (ganeshie8):

simply swap the number in your list : 1-2 2-3 3-4 4-5 5-6

MARC (marc_d):

ok..

ganeshie8 (ganeshie8):

How many outcomes are in our favor ?

MARC (marc_d):

12

ganeshie8 (ganeshie8):

I thought we had 10+4 = 14. How did you get 12 ?

MARC (marc_d):

xD Counted wrongly.. yep B=10 and A=4

ganeshie8 (ganeshie8):

Careful with the duplicates in A and B We mustn't count the same outcome twice

MARC (marc_d):

alright

ganeshie8 (ganeshie8):

A = { (1, 4), (2, 3), (3, 2), (4, 1)} B = {(2, 1), (3, 2), (4, 3), (5, 4), (6, 5), (1, 2), (2, 3), (3, 4), (4, 5), (5, 6)}

ganeshie8 (ganeshie8):

A = { (1, 4), \(\color{red}{(2, 3), (3, 2)}\), (4, 1)} B = {(2, 1), \(\color{red}{(3, 2)}\), (4, 3), (5, 4), (6, 5), (1, 2), \(\color{red}{(2, 3)}\), (3, 4), (4, 5), (5, 6)}

ganeshie8 (ganeshie8):

A has 4 elements B has 10 elements Both have 2 elements in common

MARC (marc_d):

yes

ganeshie8 (ganeshie8):

So total number of distinct elements = ?

MARC (marc_d):

12

ganeshie8 (ganeshie8):

How ?

MARC (marc_d):

14-2

ganeshie8 (ganeshie8):

12 was right, I'm just asking how you worked it

MARC (marc_d):

bcoz it has 2 common

ganeshie8 (ganeshie8):

Yes! \[n(A ~\cup~ B) = n(A) + n(B) - n(A ~\cap ~B)\]

ganeshie8 (ganeshie8):

In words : (number of elements in A or B) = (number of elements in A) + (number of elements in B) - (number of elements in both A and B)

ganeshie8 (ganeshie8):

12 is the number of outcomes in favor

ganeshie8 (ganeshie8):

rest should be easy ?

MARC (marc_d):

yes.

MARC (marc_d):

P(A)=1/3 P(B)=5/6 P(A and B)=1/6

ganeshie8 (ganeshie8):

Hey we don't have to calculate the individual probabilities

ganeshie8 (ganeshie8):

You have number of outcomes in favor (12) You knew the total number of outcomes (36) simply divide and you're done!

MARC (marc_d):

1/3 Thank you @ganeshie8 ^_^

ganeshie8 (ganeshie8):

Alternatively, you could also find the individual probabilties

ganeshie8 (ganeshie8):

A=getting whose sum is 5 B=getting their difference is 1 P(A) = ? P(B) = ?

MARC (marc_d):

\[P(A)=\frac{ 4 }{ 36 }\]

ganeshie8 (ganeshie8):

Yes

MARC (marc_d):

P(B)=10/36=5/18

MARC (marc_d):

P(A and B)=2/36

ganeshie8 (ganeshie8):

Yes! P(A or B) = P(A) + P(B) - P(A and B)

ganeshie8 (ganeshie8):

plugin the numbers and you should get the same answer

MARC (marc_d):

yep,i'm getting 1/3 by using that formula.. ^ Thanks again @ganeshie8 :)

ganeshie8 (ganeshie8):

Np :)